OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! Solve the differential equation dy dx equals the quotient of y times the cosine of x and the quantity 1 plus y squared with the initial condition y(0) = 1.

OpenStudy (aryana_maria2323):

Can you please help? @mathmale @Nnesha @Directrix

Nnesha (nnesha):

\[\large\rm \frac{ dy }{ dx }=\frac{ ycos(x) }{ (1+y^2) }\] like this or \[\large\rm \frac{ dy }{ dx }=\frac{ ycos(x) }{ (1+y)^2 }\]

Nnesha (nnesha):

?

Nnesha (nnesha):

hi ??

Nnesha (nnesha):

well i gtg so i think 2nd one is correct first isolate dy \[\large\rm 1dy=\frac{ ycos(x) }{ (1+y)^2 } dx\] and then integrate \[\large\rm \int\limits_{}^{}1dy=\int\limits_{}^{}\frac{ ycos(x) }{ (1+y)^2 } dx\] find the anti-derivative of both sides y(0)=1 represents that when x=0 y=1 so at the end substitute x and y for these numbers

OpenStudy (mathmale):

If other users have to ask you for clarification about what you meant by a verbal descrption of a math problem situation, that's a clue that you should perhaps be honing your presentation skills using math symbols instead of words.

OpenStudy (aryana_maria2323):

I am so sorry I had to help my dad with something. The first one is correct. @mathmale can you please help me?

OpenStudy (anonymous):

separate the variables and integrate \[\int\limits \frac{ 1+y^2 }{ y }dy=\int\limits \cos x dx+c\]

OpenStudy (aryana_maria2323):

Ohh okay is that is?

OpenStudy (aryana_maria2323):

@surjithayer

OpenStudy (aryana_maria2323):

@surjithayer

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ y }dy+\int\limits y~dy=\sin x+c\] \[\ln y+\frac{ y^2 }{ 2 }=\sin x+c\] when x=0,y=1 \[\ln 1=0\]

OpenStudy (anonymous):

find c and then substitute the value of c

OpenStudy (aryana_maria2323):

C=-sin(x)+ln(y)+y^2/2 is that correct? @surjithayer

OpenStudy (aryana_maria2323):

Am I right @Nnesha

Nnesha (nnesha):

that's where you need to use the given info y(0)=1 when x=0 y=1 substitute x for 0 y for 1

OpenStudy (aryana_maria2323):

okay but what part of the equation do I use. there are so many steps I am not 100% sure where to plug in 0 and 1

Nnesha (nnesha):

you're good to plugin 0 and 1 right after you find out the anti derivative

OpenStudy (aryana_maria2323):

what do I find the anti derivative of?

OpenStudy (aryana_maria2323):

@Nnesha

Nnesha (nnesha):

hmm look at surjithayer's comment that's a good way to start

Nnesha (nnesha):

to get rid of the fraction and multiplication he moved (y/1+y^2) to the left side

OpenStudy (aryana_maria2323):

C=1/2?

Nnesha (nnesha):

how did you get 1/2 ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @aryana_maria2323 C=-sin(x)+ln(y)+y^2/2 is that correct? \(\color{blue}{\text{End of Quote}}\) substitute x for 0 y=1

OpenStudy (aryana_maria2323):

ln(1)+1^2/2=sin(0)+C C=1/2

Nnesha (nnesha):

ye.

OpenStudy (aryana_maria2323):

So im correct?

Nnesha (nnesha):

yes

OpenStudy (aryana_maria2323):

Yay thank you so much.